# Old Meets New: Connecting Two Infinite Families of Congruences Modulo Powers of 5 for Generalized Frobenius Partition Functions

### v2

@@ Abstract.
In 2012 Paule and Radu proved a difficult family of congruences modulo powers of 5 for Andrews’ 2-colored generalized Frobenius partition function. The family is associated with the classical modular curve of level 20. We demonstrate the existence of a congruence family for a related generalized Frobenius partition function associated with the same curve. Unlike traditional proofs, we construct an isomorphism between this new family and the original family of congruences via a mapping on the associated rings of modular functions. The existence of this second congruence family and its relationship with the first poses significant implications for future work on congruences associated with modular curves of composite level. We show how a similar approach can be made to multiple other recent examples in the literature. We also give some important insights into the behavior of these congruence families with respect to the Atkin–Lehner involution which proved very important in Paule and Radu’s original proof.
@@ 1. Introduction
It appears that a remarkable property exists for congruence families associated with modular curves of composite level—in some sense, the same family can manifest in the Fourier coefficients associated with multiple different modular forms. We show how our approach may be applied to many other recently studied partition functions.
@@ 1.1. Background
The third author has invented some more general techniques, embodied in the localization method, in an attempt to give a unified framework to the subject of partition congruences. These techniques work in many different cases (e.g., [19], [20]), and indeed reduce to standard methods for classical congruence families. However, the congruences of Theorem 1.1 are strangely resistant to localization. As such, we cannot at present extend classical methods via a systematic approach to congruence families of a certain class which encompass Theorem 1.1 and other contemporary congruence families, some of which are yet standing conjectures. In a sense, we know that the family is correct, but we still do not fully understand why it is there, and why more systematic approaches fail to resolve it.
@@
It appears that in some strange sense Theorems 1.1 and 1.2 constitute a “single” congruence family (i.e., a specific pattern of $5$-adic convergence to 0), which somehow manifests itself in two different ways, for two different functions. This approach may have enormous implications on the future study of congruence families associated with modular curves of composite level.
@@
Finally, these families bear a strange symmetry to one another upon the application of a certain Atkin–Lehner involution to their respective functions. This symmetry gives rise to an absolutely astonishing and beautiful result which explains a curious coefficient behavior that was discovered by Paule and Radu in their original work [15, Section 6], and used to assist in the construction of their proof of Theorem 1.1.
@@
The second example involves a pair of altogether different functions sudied by Connor Morrow in collaboration with the first author [11]. One is a congruence family exhibited by a function related to the crank parity function originally proved by Choi, Kang, and Lovejoy in [7]. The other is a family exhibited by a function closely related to the 1-shell TSPP and $\overline{\mathcal{EO}}(n)$ functions studied in [2], [5], and [12]. Moreover, we note that yet more extremely recent work in this subject has been done by the first author in concert with Dandan Chen and Rong Chen [4].
Still another striking example involves a remarkable pair of infinite internal congruence families exhibited by coefficients related to quotients of Ramanujan’s theta functions, studied by Chern and Tang [6].
@@
In this work we presume an understanding of the theory of modular functions and its associated modular curves. See [19, Section 2]. A useful classical reference is [14], and a more modern reference is [8, Chapters 2-3].
@@ 2. Frobenius Partition Functions
The generally accepted approach to proving a congruence family corresponding to the coefficients of a given modular form is to construct a sequence $\mathcal{L}:=(L_{\alpha})_{\alpha\geq 1}$ of functions on an associated manifold—which for our purposes will always be the classical modular curve $\mathrm{X}_{0}(N)$ for some positive integer $N$. Indeed, except for Section 5, we will always work with $\mathrm{X}_{0}(20)$. Each $L_{\alpha}$ enumerates a given case of the congruence family, and each function is related to its predecessor via a certain linear operator. The difficulty then is to express each $L_{\alpha}$ in terms of some convenient reference functions.
@@
is the standard $U_{5}$ operator (see [14, Chapter 8] for a list of the standard properties of $U_{5}$), and
@@
We instantly recognize divisibility of the function by 5 in this identity. Similar identities can easily be computed for larger $\alpha$, and it can be shown with relative ease that
@@
The possibility of actually proving the whole family by manipulation of these rational polynomials is naturally very alluring. Such a class of proofs does exist, and has been exhibited for other difficult congruence families in previous papers.
This technique fails spectacularly in the cases of the families in Theorems 1.1 and 1.2. It is certainly true that manipulation of these rational polynomial identities allows one to confirm as many cases of either theorem as we like. However, for extremely deep and technical reasons, it is not currently possible to properly complete the induction along the lines of localization.
@@
This is the Hauptmodul for $\mathrm{X}_{0}(5)$, a curve of genus 0 and cusp count 2. Indeed, $t$ is the critical function for proving Ramanujan’s congruence family for $p(n)$ modulo powers of 5 (see [22] or [14, Chapters 7-8]). Being modular over $\Gamma_{0}(5)$, it is clearly modular over $\Gamma_{0}(20)$. It does not live at the cusp $[0]$, but we can perform the same trick we did with $L_{1}^{(1)}$—we can represent $t$ as a rational polynomial. Indeed, we have
$t=\frac{x(1+4x)^{2}}{1+5x}$ (3.14)
(this is proved by a straightforward procedure in our Mathematica supplement, which may be found online in the url at the end of Section 1). Notice that if we multiply through by $1+5x$, substitute $z=1+5x$, and clear fractions, we then have
@@
However, the existence of such a ring mapping system as $\sigma^{(\beta)}$ strongly suggests that there exists some explicit means of transforming the functions corresponding to $c\psi_{2,0}(n)$ into those corresponding to $c\psi_{2,1}(n)$. Our first instinct on the matter was to attempt a connection between $\mathrm{C}\Psi_{2,\beta}$, $\mathrm{C}\Psi_{2,1-\beta}$ via an Atkin–Lehner involution.
@@
which of course will reduce via the associated transformation to
@@
Notice that by this transformation, we immediately have
@@
Thus $\sigma^{(1)}$ equates to sending $f(\tau)$ to $f(\gamma\tau)$, and we can define
@@
Supposing that for some $\alpha\geq 2$ we have $L_{\alpha-1}^{(0)}=L_{\alpha-1}^{(1)}(\gamma\tau)$. Again allowing $\alpha\equiv a\pmod{2}$ for $a\in\{0,1\}$, we have
@@ 4.1. Congruences Related to Other Frobenius Partitions
Thus the equivalences that we are constructing between various different partition functions imply equivalences of nonexistence, as well as existence, of divisibility properties of interest.
@@ 4.3. Internal Congruences Related to Ramanujan’s Theta Functions
Yet another extremely fascinating example of such equivalences of congruences comes from an elegant study [6] by Chern and Tang of two families of internal congruences modulo powers of 3 for certain quotients of Ramanujan’s theta functions $\varphi(q)$, $\psi(q)$. Let
@@
Just as in the previous examples of this section, the phenomenon of matching coefficients for the polynomial representations of $\hat{\Phi}_{\alpha}$ and $\hat{\Psi}_{\alpha}$ strongly suggests that an approach analogous to our work on $c\psi_{2,\beta}(n)$ should be applicable. Indeed, let us take the Atkin–Lehner involution
@@ 6. The Atkin–Lehner Involution
We close our report with a discussion of the striking dual relationship our congruence pair has with respect to Ramanujan’s classic $\varphi(q)$ function via a certain Atkin–Lehner involution. This partly answers (or at least extends) some of the questions which Paule and Radu posed in their proof of Theorem 1.1. Recall the Atkin–Lehner involution from Section 3.1 which arises from the action of the matrix
@@
We thus have a means of studying the direct effect of $W$ on our function spaces. Now, Paule and Radu discovered that
@@
With these, we have the following result, beautiful and strange:
@@ 8. Acknowledgments
The third author was funded in whole by the Austrian Science Fund (FWF) Principal Investigator Project 10.55776/PAT6428623, “Towards a Unified Theory of Partition Congruences.”

### v3

@@ Abstract.
In 2012 Paule and Radu proved a difficult family of congruences modulo powers of 5 for Andrews’ 2-colored generalized Frobenius partition function. The family is associated with the classical modular curve of level 20. We demonstrate the existence of a congruence family for a related generalized Frobenius partition function associated with the same curve. We construct an isomorphism between this new family and the original family of congruences via a mapping on the associated rings of modular functions. The pairing of the congruence families provides a new strategy for future work on congruences associated with modular curves of composite level. We show how a similar approach can be made to multiple other recent examples in the literature. We also give some important insights into the behavior of these congruence families with respect to the Atkin–Lehner involution which proved very important in Paule and Radu’s original proof.
@@ 1. Introduction
It appears that a remarkable property exists for congruence families associated with modular curves of composite level—in some sense, the same family can manifest in the Fourier coefficients associated with multiple modular forms. We show how our approach may be applied to many other recently studied partition functions.
@@ 1.1. Background
The third author has invented some more general techniques, embodied in the localization method, in an attempt to give a unified framework to the subject of partition congruences. These techniques revolve around manipulation of certain rational polynomial sequences of modular functions to extract congruence properties. This approach works in many different cases (e.g., [20], [21], [4]), and indeed it reduces to standard methods for classical congruence families; however, this approach fails to resolve the congruences of Theorem 1.1. As such, we cannot at present extend classical methods via a systematic approach to congruence families of a certain class which encompass Theorem 1.1 and other contemporary congruence families, some of which are yet standing conjectures. We know that the family has been proved, but we still do not fully understand why it is there, and why more systematic approaches fail to resolve it.
@@
It appears that in some strange sense Theorems 1.1 and 1.2 constitute a “single” congruence family (i.e., a specific pattern of $5$-adic convergence to 0), which somehow manifests itself in two different ways, for two different functions. This approach provides a new strategy for the future study of congruence families associated with modular curves of composite level.
@@
Finally, these families bear a strange symmetry to one another upon the application of a certain Atkin–Lehner involution to their respective functions. This symmetry gives rise to an absolutely astonishing and beautiful result which explains a curious coefficient behavior that was discovered by Paule and Radu in their original work [16, Section 6], and used to assist in the construction of their proof of Theorem 1.1.
@@
The second example involves a pair of altogether different functions sudied by Connor Morrow in collaboration with the first author [12]. One is a congruence family exhibited by a function related to the crank parity function originally proved by Choi, Kang, and Lovejoy in [8]. The other is a family exhibited by a function closely related to the 1-shell TSPP and $\overline{\mathcal{EO}}(n)$ functions studied in [2], [6], and [13]. Moreover, we note that yet more recent work in this subject has been done by the first author in concert with Dandan Chen and Rong Chen [5].
Still another example involves a pair of infinite internal congruence families exhibited by coefficients related to quotients of Ramanujan’s theta functions, studied by Chern and Tang [7].
@@
In this work we presume an understanding of the theory of modular functions and its associated modular curves. See [20, Section 2]. A useful classical reference is [15], and a more modern reference is [9, Chapters 2-3].
@@ 2. Frobenius Partition Functions
A common approach to proving a congruence family corresponding to the coefficients of a given modular form is to construct a sequence $\mathcal{L}:=(L_{\alpha})_{\alpha\geq 1}$ of functions on an associated manifold—which for our purposes will always be the classical modular curve $\mathrm{X}_{0}(N)$ for some positive integer $N$. Indeed, except for Section 5, we will always work with $\mathrm{X}_{0}(20)$. Each $L_{\alpha}$ enumerates a given case of the congruence family, and each function is related to its predecessor via a certain linear operator. The difficulty then is to express each $L_{\alpha}$ in terms of some convenient reference functions.
@@
is the standard $U_{5}$ operator (see [15, Chapter 8] for a list of the standard properties of $U_{5}$), and
@@
We recognize divisibility of the function by 5 in this identity. Similar identities can easily be computed for larger $\alpha$, and it can be shown that
@@
The possibility of actually proving the whole family of congruences by manipulation of these rational polynomials is naturally very alluring. Such a class of proofs does exist, and has been exhibited for other difficult congruence families in previous papers, e.g., [20], [21], [4].
This technique fails spectacularly in the cases of the families in Theorems 1.1 and 1.2. It is certainly true that manipulation of these rational polynomial identities allows one to confirm as many cases of either theorem as we like. However, it is not currently possible to complete the induction along these lines owing to our current inability to properly stabilize the associated congruence ideal sequence (see [4, Section 6] for an explicit description of this approach).
@@
This is the Hauptmodul for $\mathrm{X}_{0}(5)$, a curve of genus 0 and cusp count 2. Indeed, $t$ is the critical function for proving Ramanujan’s congruence family for $p(n)$ modulo powers of 5 (see [23] or [15, Chapters 7-8]). Being modular over $\Gamma_{0}(5)$, it is clearly modular over $\Gamma_{0}(20)$. It does not live at the cusp $[0]$, but we can perform the same trick we did with $L_{1}^{(1)}$—we can represent $t$ as a rational polynomial. Indeed, we have
$t=\frac{x(1+4x)^{2}}{1+5x}.$ (3.14)
Notice that if we multiply through by $1+5x$, substitute $z=1+5x$, and clear fractions, we then have
@@
However, the existence of such a ring mapping system as $\sigma^{(\beta)}$ suggests that there exists some explicit means of transforming the functions corresponding to $c\psi_{2,0}(n)$ into those corresponding to $c\psi_{2,1}(n)$. Our first instinct on the matter was to attempt a connection between $\mathrm{C}\Psi_{2,\beta}$, $\mathrm{C}\Psi_{2,1-\beta}$ via an Atkin–Lehner involution.
@@
which will reduce via the associated transformation to
@@
Notice that by this transformation, we have
@@
Thus $\sigma^{(1)}$ sends $f(\tau)$ to $f(\gamma\tau)$, and we can define
@@
Supposing that for some $\alpha\geq 2$ we have $L_{\alpha-1}^{(0)}=L_{\alpha-1}^{(1)}(\gamma\tau)$. Again allowing $\alpha\equiv a\pmod{2}$ for $a\in\{0,1\}$,
@@ 4.1. Congruences Related to Other Frobenius Partitions
Thus the equivalences that we are constructing between various partition functions imply equivalences of nonexistence, as well as existence, of divisibility properties of interest.
@@ 4.3. Internal Congruences Related to Ramanujan’s Theta Functions
Yet another example of such equivalences of congruences comes from an elegant study [7] by Chern and Tang of two families of internal congruences modulo powers of 3 for certain quotients of Ramanujan’s theta functions $\varphi(q)$, $\psi(q)$. Let
@@
As in the previous examples of this section, the phenomenon of matching coefficients for the polynomial representations of $\hat{\Phi}_{\alpha}$ and $\hat{\Psi}_{\alpha}$ strongly suggests that an approach analogous to our work on $c\psi_{2,\beta}(n)$ should be applicable. Indeed, let us take the Atkin–Lehner involution
@@ 6. The Atkin–Lehner Involution
We close our report with a discussion of the dual relationship our congruence pair has with respect to Ramanujan’s classic $\varphi(q)$ function via a certain Atkin–Lehner involution. This partly answers (or at least extends) some of the questions which Paule and Radu posed in their proof of Theorem 1.1. Recall the Atkin–Lehner involution from Section 3.1 which arises from the action of the matrix
@@
We thus have a means of studying the effect of $W$ on our function spaces. Now, Paule and Radu discovered that
@@
With these, we have the following result:
@@ 8. Acknowledgments
This research was funded in whole or in part by the Austrian Science Fund (FWF) Principal Investigator Project 10.55776/PAT6428623, “Towards a Unified Theory of Partition Congruences.” For open access purposes, the authors have applied a CC BY public copyright license to any author-accepted manuscript version arising from this submission.
In particular, the third author, Nicolas Allen Smoot, is the Principal Investigator of the aforementioned FWF Project, and he would like to thank the Austrian Government and People for their generous support.
All three authors would like to thank the anonymous referees, who provided careful and constructive advice which substantially improved this manuscript.